hi.. i need help with writing and solving an equation in standard form. for instance.. y- intercept=8 and the slope is 3. does that mean the equation in standard form is 3x+y=-8?? I also need help with.. if the slope = 3/4 and y-intercept= -2 what's the equation in standard form and how can you get 3/4 to not be a fraction or negative? same with slope= 3/5 and passes through 0, -6.. i think that means the y-int. is -6 but how do you graph slopes that are fractions and write them in an equation. sorry it's such a long question but please help me.. thanks!!
Answer by venugopalramana(3286)

hi.. i need help with writing and solving an equation in standard form. for instance.. y- intercept=8 and the slope is 3. does that mean the equation in standard form is 3x+y=-8?? I also need help with.. if the slope = 3/4 and y-intercept= -2 what's the equation in standard form and how can you get 3/4 to not be a fraction or negative? same with slope= 3/5 and passes through 0, -6.. i think that means the y-int. is -6 but how do you graph slopes that are fractions and write them in an equation. sorry it's such a long question but please help me.. thanks!!
The slope intercept form of eqn.of a st.line is y=mx+c..where m is the slope and c is the y intercept.so your answer should be
..y=3x+8 and not y+3x=8
now slope=3/4.. we are given that y intercept is –2.. so the eqn.is..
..y=(3/4)x-2….to remove the fraction multiply both sides of the eqn.with 4 we get
4y=(3*4/4)x-2*4…or …4y=3x-8…if you so desire you put all of them on one side as
4y-3x+8=0
now slope is 3/5 and passes through 0, -6.. i think that means the y-int. is -6 YES CORRECT YOU ARE ..so proceed as above to get the eqn.and remove fraction in slope as is convenient.the resulting eqn.you should be able to graph easily by giving different values for x say 0,1,2,3..etc and finding corresponding values of y and plotting.
************